### v1

# MAESTRO: Multi-Agent Environment Shaping through Task and Reward Optimization

Abstract
Cooperative Multi-Agent Reinforcement Learning (MARL) is fundamentally constrained by two design bottlenecks: the manual engineering of dense reward functions and the difficulty of constructing curricula that prevent local optima in high-dimensional state spaces. Existing solutions typically rely on static, hand-crafted heuristics that struggle to generalize. While Large Language Models (LLMs) have recently been applied to decision-making, most approaches utilize the “LLM-as-Agent” paradigm—placing the model directly in the control loop—which incurs prohibitive inference latency and token costs unsuitable for real-time systems.
In this work, we propose MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), a framework that shifts the role of the LLM from real-time agent to high-level architect. MAESTRO treats the training process as a dual-loop optimization problem where the LLM dynamically designs both the task and the solving guidance. The system integrates two generative components: (i) a semantic curriculum generator that synthesizes diverse, context-rich scenarios based on agent performance, and (ii) an automated reward synthesizer that writes executable Python reward functions to guide agents through the evolving curriculum. These components optimize the training of a standard off-policy MARL backbone (MADDPG) without requiring LLM inference during execution.
[…]
Our code is available at: https://github.com/BrianWu1010/MAESTRO.git.
1 Introduction
Cooperative Multi-Agent Reinforcement Learning (MARL) holds immense promise for solving complex decentralized control problems, from warehouse logistics to sensor network coordination. However, scaling MARL to realistic, non-stationary environments is persistently hindered by two fundamental design bottlenecks: curriculum design and reward engineering.
The first bottleneck, exploration, is often addressed by **Curriculum Learning (CL)**, which gradually increases task difficulty. However, constructing effective curricula requires deep domain expertise to ensure tasks are meaningfully diverse and effective [narvekar2020curriculum]. The second bottleneck concerns reward specification: the efficacy of MARL is brittle with respect to reward specification. Standard dense rewards are often misaligned with global objectives [zang2020metalight], while sparse global rewards fail to provide sufficient learning signal. Handcrafting reward functions that generalize across dynamic regimes remains a labor-intensive and error-prone process.
Large Language Models (LLMs) offer a potential remedy by encoding rich semantic priors that can automate these design processes. However, the prevailing integration paradigm—LLM-as-Agent—typically inserts the model directly into the control loop to select actions [lai2023llm, da2023llm, zhang2023trafficgpt]. While this yields strong zero-shot generalization, it incurs prohibitive inference latency and token costs, rendering it impractical for real-time control systems requiring sub-second decision making. Furthermore, this approach underutilizes the LLM’s capacity for high-level reasoning, relegating it to per-step tactical decisions.
Our Approach: MAESTRO.
In this work, we propose a paradigm shift from LLM-as-Agent to LLM-as-Architect. We introduce MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), a generative meta-learning framework that uses the LLM offline to shape the training process of a conventional MARL backbone. We treat the training loop as a dual-optimization problem in which the LLM acts as a high-level designer. MAESTRO integrates two generative components: (i) a Semantic Curriculum Generator that synthesizes diverse, performance-driven task contexts, and (ii) an Automated Reward Synthesizer that produces executable Python reward functions dynamically tailored to the current curriculum stage.
By distilling semantic knowledge into executable training scaffolds (tasks and rewards), MAESTRO guides a standard MADDPG policy, enabling the final agents to retain the lightweight, low-latency characteristics required for deployment while benefiting from the broad reasoning capabilities of foundation models.
Figure 1: The MAESTRO framework workflow. The system operates as a dual-loop process: The LLM (Architect) designs the learning environment by generating curriculum contexts and synthesizing executable Python reward functions. The MARL agents (Learners) train within this shaped environment. Performance feedback triggers the Architect to adapt the curriculum and regenerate rewards, isolating the expensive LLM inference from the real-time execution loop.
Contributions.
We validate MAESTRO on the challenging domain of Urban Traffic Signal Control (TSC), a prominent benchmark for MARL coordination [wei2019colight, zhang2019cityflow]. TSC specifically serves as a rigorous testbed for non-stationarity and coordination. Our contributions are:
  • A Generative Environment Shaping Framework. We formally propose MAESTRO, a method for contextual MARL that automates both task generation and reward synthesis, shifting the LLM’s role to an offline supervisor.
  • Controlled Ablation of Generative Components. We conduct a rigorous ablation comparing three configurations: a baseline LLM-driven curriculum with template rewards (A2), an alternative curriculum mode (A8), and the full MAESTRO system with LLM-generated reward shaping (A7). This design explicitly isolates the value of generative reward shaping from curriculum learning.
[…]
2 Related Work
2.1 MARL for Decentralized Control
Multi-Agent Reinforcement Learning (MARL) is widely applied to cooperative control problems where agents share a global goal but make local decisions based on partial observations. TSC, a challenging instance of a Dec-POMDP, popularized the Centralized Training with Decentralized Execution (CTDE) paradigm [lowe2017maddpg], where agents learn a joint value function during training but act independently in deployment. Early applications of CTDE to traffic, such as CoLight [wei2019colight] and MPLight [chen2020mplight], focused on modeling inter-agent dependencies using mechanisms like graph attention and phase-pressure signals. These approaches assume a static, manually designed reward function and a fixed environment setup. Our work maintains the stability and real-time capability of the CTDE backbone (MADDPG) but allows the reward function itself and the training environment to evolve dynamically via LLM-generated shaping signals, addressing a critical generalization gap.
2.2 Automated Curriculum and Environment Generation
The difficulty of exploration in high-dimensional MARL motivates Automatic Curriculum Learning (ACL), which autonomously sequences tasks to maximize agent learning progress [portelas2020automatic]. Methods like POET [wang2019paired] and ALP-GMM [portelas2020automatic] demonstrate the power of generating diverse environments to force agents to acquire generalizable skills. In the context of TSC, curriculum generation has traditionally been limited to heuristic, numeric adjustments of traffic parameters, such as flow rates and phase durations [zang2020metalight]. These numeric methods lack the expressiveness to generate complex, semantically meaningful scenarios (e.g., “a peak-hour commuter pattern with a major bottleneck on the eastern artery”). Our approach moves beyond numeric scaling to a language-informed ACL, where an LLM generates semantically rich traffic contexts conditioned on a scalar difficulty signal, significantly expanding the scope of environment shaping.
2.3 LLMs for Reward and Policy Generation
The integration of Large Language Models into RL systems is broadly categorized into two paradigms:
LLM-as-Agent.
In this setting, the LLM directly selects actions or high-level plans, often via chain-of-thought reasoning. TrafficGPT [zhang2023trafficgpt] and LLM-TSCS [lai2023llm] exemplify this for traffic control. While providing strong zero-shot behavior, these methods are bottlenecked by the LLM’s high inference latency and token cost, making them ill-suited for real-time control applications with strict timing constraints.
LLM-as-Teacher (Offline Guidance).
This paradigm leverages the LLM’s reasoning capabilities during the training phase to shape the learning process. Recent work like Eureka [ma2023eureka] and Code-as-Policies [liang2023code] demonstrated that LLMs can generate effective policy code or, critically, executable reward functions for single-agent or robotic skill acquisition. These reward functions act as powerful, automatically generated priors.
MAESTRO builds upon and substantially extends the LLM-as-Teacher direction for the cooperative MARL domain. We introduce two key novelties: (i) the generative process produces executable Python reward functions for a decentralized, cooperative MARL system, which is significantly more complex than the single-agent setting; and (ii) we explicitly couple reward generation with a performance-driven, LLM-generated curriculum, ensuring the reward logic evolves in synergy with the increasing task difficulty. This coupling is essential for the stability and generalization of MARL policies in non-stationary environments.
3 Preliminaries
3.1 Contextual MARL Formulation and Objectives
We model multi-intersection TSC as a contextual Decentralized Partially Observable Markov Decision Process (Dec-POMDP), augmented with the MAESTRO training framework. The environment is defined as
$\mathcal{M}_{C}=\langle\mathcal{C},\mathcal{I},\mathcal{S},\{A_{i}\}_{i\in\mathcal{I}},\{O_{i}\}_{i\in\mathcal{I}},\mathcal{T}_{C},\mathcal{R}_{C},\gamma\rangle,$
where:
  • $\mathcal{C}$ denotes the curriculum context space (generated by the LLM Architect), which parameterizes traffic flows, route patterns, and demand distributions.
  • $\mathcal{I}=\{1,\dots,N\}$ is the set of traffic-signal agents (intersections).
  • $\mathcal{S}$ is the global state; $O_{i}$ and $A_{i}$ are the observation and action spaces for agent $i$.
  • $\mathcal{T}_{C}$ is the context-dependent transition function.
  • $\mathcal{R}_{C}$ is the context-dependent reward function (synthesized by the LLM Architect), and $\gamma$ is the discount factor.
At the start of each training stage, a context $C\in\mathcal{C}$ is sampled and the LLM Architect produces the corresponding reward function $\mathcal{R}_{C}$. Agents then interact with the environment under the joint policy $\pi_{C}$, generating a trajectory $\tau$. The objective is to maximize the expected cumulative return over all curriculum contexts:
$\max_{\pi}\;\mathbb{E}_{C\sim p(\mathcal{C}),\;\tau\sim\mathcal{T}_{C}(\cdot\mid\pi_{C})}\left[\sum_{t=0}^{T}\gamma^{t}\mathcal{R}_{C}(s_{t},a_{t})\right].$ (1)
We adopt MADDPG [lowe2017maddpg] as the MARL Learner backbone within the Centralized Training with Decentralized Execution (CTDE) paradigm. This design preserves stability during cooperative learning while maintaining efficient, decentralized execution at test time—cleanly separating the LLM’s offline design role from the agents’ real-time control.
3.2 MARL Learner Specification
Simulation Setup.
Experiments use the CityFlow microscopic simulator [zhang2019cityflow] on the real-world Hangzhou (HZ) network with $|\mathcal{I}|=16$ signalized intersections. Each episode spans 360 timesteps (1s per step). The curriculum context $C$ configures dynamic elements such as route distributions, vehicle arrival patterns, and vehicle parameters. Full parameter ranges are provided in the appendix.
Observations and Actions.
Each agent $i$ receives a structured observation vector $o_{i}^{t}$ comprising local queue lengths, inflow/outflow rates, current phase, elapsed phase duration, lane occupancy, and short-horizon traffic history. This feature design follows established MARL–TSC practice and is detailed in the appendix. Agent actions correspond to selecting one of the $|P_{i}|=4$ feasible signal phases at each timestep, consistent with intersection geometry.
Traffic Metrics.
Performance is evaluated using standard TSC metrics: mean travel time, delay, queue length, and throughput. In MAESTRO, these metrics serve two purposes: (i) they drive difficulty adjustment for the curriculum generator, and (ii) they are combined into the composite reward signal $\mathcal{R}_{C}$ used during agent training.
4 Methodology
This section introduces MAESTRO and the three evaluated configurations. Figure 1 illustrates the overall workflow, and Figure 2 shows the conceptual foundation.
Figure 2: Conceptual foundation of the MAESTRO architecture. The LLM acts as the Architect, generating task-specific, executable components (e.g., reward functions and curriculum contexts) that are injected into the MARL training loop. MAESTRO focuses on LLM-generated reward shaping coupled with an adaptive curriculum, while keeping the underlying MARL Learner backbone (MADDPG) unchanged for deployment safety and real-time efficiency.
4.1 Overview of MAESTRO Configurations
We evaluate three configurations that differ only in how they utilize the LLM for generative curriculum design and reward shaping. All other components (network, MADDPG architecture, and hyperparameters) remain identical, enabling a clean controlled ablation:
Config Curriculum Mode Reward Source LLM Reward Shaping?
A2 llm_adaptive Template-based (Environment) No
A8 llm Template-based (Environment) No
A7 llm_adaptive Environment + LLM-shaped Yes
All configurations use LLM assistance to generate curriculum contexts, but only A7 (the full MAESTRO system) incorporates the LLM-generated auxiliary reward. The ablation focuses on isolating the contribution of generative reward shaping within an already LLM-assisted curriculum.
4.2 LLM-Driven Adaptive Curriculum
The adaptive curriculum maintains a scalar difficulty $d\in[0.3,1.0]$ and updates it based on a rolling window of five episodes using the current policy’s cumulative return. Difficulty adjustment follows:
  • If the average return exceeds $1.05\times 165.0$ (173.25) for two consecutive windows, increase $d$ by $5\%$.
  • If the return drops below $0.95\times 165.0$ (156.75), decrease $d$ by $10\%$.
  • If $d$ remains unchanged for 15 episodes, apply a forced $5\%$ increase (plateau escape).
Difficulty is clipped to $[0.3,1.0]$ to avoid trivial (empty-traffic) or over-saturated scenarios.
Under llm_adaptive (A2, A7), GPT-4o-mini is prompted with the current difficulty $d$ and summary statistics to generate a semantically rich traffic context (e.g., flow patterns, demand asymmetry, vehicle parameters), which is then converted into a CityFlow configuration. Under llm (A8), a less constrained prompting and regeneration strategy is used for comparison, while applying the same difficulty-update rules. Pilot studies verified that increasing $d$ results in monotonic increases in fundamental difficulty metrics (queue length, delay) under a fixed-time policy.
4.3 LLM-Generated Reward Shaping (MAESTRO – A7)
The full MAESTRO configuration (A7) extends the A2 curriculum baseline by introducing LLM-generated auxiliary rewards that evolve with the curriculum.
Reward Generation and Validation.
At regeneration events (triggered by $\Delta d\geq 0.15$, 20 episodes since the last generation, or a stagnation-induced difficulty increase), GPT-4o-mini is prompted with the current difficulty, aggregated statistics, and a description of the learning goal. The LLM generates a short Python reward function $r_{\mathrm{LLM},t}$. A three-stage validation pipeline ensures safety:
  • Syntax Check: abstract syntax tree validation,
  • Sandboxed Execution: sample rollouts to confirm runtime behavior,
  • Safety Check: range bounding and finite-value verification.
If three consecutive generations fail, the system reverts to environment-only rewards (behaving as A2) until the next regeneration trigger.
Scaling and Difficulty-Dependent Weighting.
Because LLM-generated rewards can differ in scale, we apply a fixed scaling factor of $170.0$ to match typical environment reward magnitudes. A difficulty-dependent weight
$w(d)=w_{\min}+\frac{d-d_{\min}}{d_{\max}-d_{\min}}(w_{\max}-w_{\min}),$
with $w_{\min}=0.1$, $w_{\max}=0.5$, $d_{\min}=0.3$, and $d_{\max}=1.0$, controls the auxiliary contribution. This ensures low influence at easy difficulties and stronger guidance in challenging regimes.
Combined Reward.
The per-timestep reward used for MADDPG updates is:
$r_{t}=\begin{cases}r_{\mathrm{env},t},&\text{A2, A8 (Curriculum Baselines)},\\[3.0pt] (1-w(d))\,r_{\mathrm{env},t}+w(d)\cdot 170.0\cdot r_{\mathrm{LLM},t},&\text{A7 (Full MAESTRO)}.\end{cases}$
Here, $r_{\mathrm{env},t}$ is the CityFlow reward and $r_{\mathrm{LLM},t}$ is the auxiliary shaping signal generated by the Architect.
5 Experimental Design and Evaluation Protocol
We evaluate the MAESTRO framework by quantifying the contribution of LLM-generated reward shaping and adaptive curriculum design to the performance and robustness of cooperative MARL. Our goals are twofold: (1) isolate the effect of LLM-generated reward functions relative to template-based rewards under an adaptive curriculum, and (2) demonstrate the stability and learning efficiency of the full MAESTRO system in a complex, non-stationary environment.
Experimental Conditions.
We evaluate three configurations forming a controlled ablation (detailed in §4):
  • A2 (Curriculum Baseline): Uses the llm_adaptive curriculum with a fixed template-based environment reward.
  • A8 (Alternative Curriculum): Uses an alternative llm curriculum implementation and the same template-based reward.
  • A7 (Full MAESTRO): Uses the A2 curriculum but replaces template rewards with dynamic LLM-generated reward functions, scaled by 170.0 and weighted by a difficulty-adaptive factor $w\in[0.1,0.5]$.
All experiments use identical hyperparameters, network architecture, agent design, and simulation settings.
Key Comparisons.
This design supports two central comparisons: (1) A2 vs. A7 isolates the contribution of LLM-generated reward shaping within the same curriculum framework; (2) A2 vs. A8 evaluates two LLM curriculum strategies while keeping the reward design fixed. We train all three configurations for 200 episodes using four seeds, prioritizing deeper runs to strengthen statistical power and attribute performance differences to specific LLM components.
5.1 Environment and Curriculum Setting
All experiments use the Hangzhou (HZ) traffic network with 16 signalized intersections (§3.2). Each episode spans 360 steps. We report standard traffic-signal-control metrics: mean delay, queue length, throughput, and cumulative return.
The adaptive curriculum maintains a scalar difficulty $d$ updated from a rolling five-episode window. When the windowed return exceeds 173.25 (105% of a target 165.0) for two consecutive windows, $d$ increases by 5%. When it falls below 156.75 (95% of target), $d$ decreases by 10%. If unchanged for 15 episodes, $d$ increases by 5% to escape stagnation. The difficulty is clipped to $[0.3,1.0]$. Difficulty monotonicity with respect to task hardness (queue length and delay) is empirically verified under a fixed-time control policy. Because difficulty updates are performance-driven, curriculum trajectories differ by seed.
In the full MAESTRO system (A7), difficulty changes and stagnation events also trigger LLM reward regeneration (§4.3), whereas A2 and A8 apply identical curriculum logic but omit reward regeneration.
5.2 Metrics
We evaluate results along three dimensions:
  • Traffic Efficiency and Final Performance: Episode return, mean delay, queue length, and throughput measured over episodes 180–199.
  • Learning Efficiency and Stability: Episodes required to reach the target performance (165.0), time to reach 90% of final return, and stability measured by variance and standard deviation of returns in episodes 180–199, supplemented with a risk-adjusted metric (Sharpe Ratio).
  • Curriculum Behavior: Per-seed difficulty trajectories, counts of difficulty adjustments, and final difficulty reached.
5.3 Experimental Protocol
Each configuration (A2, A8, A7) is trained with four seeds (200, 300, 400, 500). Training runs for 200 episodes (72,000 environment steps). Evaluation episodes occur every two training episodes, with deterministic policies (no exploration noise) to measure true policy quality. Hyperparameters follow standard MADDPG practice, and all simulation parameters (network topology, demand, and CityFlow settings) remain fixed.
Reproducibility is enforced by seeding PyTorch, NumPy, Python RNG, and CuDNN. In A7, LLM-generated reward code is processed through a deterministic validation and post-processing pipeline, ensuring that identical seeds lead to identical code acceptance decisions.
Statistical analyses operate on seed-level aggregates. For each configuration and seed, we compute mean return over episodes 180–199 as the summary of final performance. Pairwise comparisons use the Mann–Whitney U test ($\alpha=0.05$, Bonferroni corrected). Effect sizes use Cohen’s $d$, and uncertainty is expressed via 95% bootstrap confidence intervals computed from 10,000 resamples. We avoid pooling episodes across seeds to preserve independence and capture between-seed variance.
5.4 Visualization Outputs
Results are visualized through four key figures: (1) learning curves showing episode return over training with 95% confidence bands (Fig. 3); (2) box plots with per-seed markers summarizing final performance (Fig. 4); (3) ablation study comparing component contributions across A2, A8, and A7 (Fig. 5); and (4) stability analysis via coefficient of variation comparison (Fig. 6).
6 Results
We evaluate MAESTRO through a comprehensive analysis of learning dynamics, final performance, traffic efficiency, and training stability across the three configurations: A2 (Curriculum Baseline), A7 (Full MAESTRO), and A8 (Alternative Curriculum). Figure 3 illustrates the learning curves, Figure 4 presents final performance distributions, and Table 2 summarizes the quantitative outcomes.
6.1 Traffic Efficiency Improvements
Both MAESTRO-enhanced configurations (A7, A8) outperform the baseline A2 in late-training traffic efficiency metrics. Table 1 reports representative results for seed 200. A8 yields the largest improvements, reducing mean delay by 27.8% and increasing throughput by 9.2%. A7 also demonstrates substantial gains, decreasing delay by 19.9%. These improvements confirm that both curriculum modulation (A8) and reward shaping (A7) provide meaningful benefits in traffic throughput and congestion mitigation.
Table 1: Traffic efficiency metrics during late-training episodes (seed 200). Higher throughput and lower temporal metrics indicate better performance. Percentages denote relative changes compared to A2.
Condition Throughput Avg Travel Avg Delay Avg Wait
(veh/ep) Time (s) (s) Time (s)
A2 (LLM Curr. + Template) 1973 793 472 502
A7 (Full MAESTRO) 2090 (+5.9%) 739 ($6.9\%$) 378 ($19.9\%$) 422 ($15.8\%$)
A8 (Alt. Curr. + Template) 2155 (+9.2%) 708 ($10.8\%$) 341 ($27.8\%$) 383 ($23.6\%$)
Figure 3: Training performance across 200 episodes. Solid lines denote mean return; shaded regions indicate 95% confidence intervals. A7 shows the narrowest confidence band, indicating the highest training stability.
6.2 Final Performance and Stability Analysis
Table 2 reports final performance statistics (episodes 180–199). A8 achieves the highest mean return (160.35), but with markedly higher variance (CV 2.55%), consistent with the wide spread seen in Figure 4. In contrast, A7 attains moderate performance gains (158.87) with exceptionally low variance (CV 0.65%), reflecting highly consistent learning dynamics across seeds.
Table 2: Ablation summary of episode returns over episodes 180–199 (higher is better). CV: coefficient of variation. Sharpe Ratio (SR) = (mean – A2 mean)/std.
Condition Curric. Reward Mean $\pm$ SD CV (%) SR Best Worst
A2 (Baseline) llm_adaptive Template $157.04\pm 2.66$ 1.69 159.63 152.82
A7 (Full MAESTRO) llm_adaptive LLM-shaped $158.87\pm 1.04$ 0.65 1.76 160.15 157.76
A8 (Alt. Curric.) llm Template $160.35\pm 4.08$ 2.55 0.81 166.05 154.77
A key observation is that MAESTRO’s reward shaping mechanism (A7) yields robust, risk-adjusted improvements. While A8 provides stronger peak performance, its variability makes it less reliable. A7 achieves a Sharpe Ratio of 1.76 (more than twice A8’s 0.81), driven by consistent performance across all seeds.
Figure 4: Distribution of final episode returns (episodes 180–199). A7 exhibits tight clustering around the mean, whereas A8 shows large variability, consistent with Table 2.
6.3 Ablation Insights
Curriculum Mode (A2 vs. A8).
Switching from llm_adaptive (A2) to llm (A8) yields a notable increase in mean return, confirming that more expressive curriculum descriptions can produce harder and more informative training samples. However, the substantially higher variance indicates that unconstrained curriculum generation can produce unstable learning trajectories.
Reward Shaping (A2 vs. A7).
Introducing LLM-generated reward functions yields consistent gains under an identical curriculum schedule. The dramatically reduced variance of A7 supports the interpretation that learned reward functions act as structured priors, damping out high-variance exploration and stabilizing policy updates.
Stability–Performance Trade-off (A7 vs. A8).
A7 and A8 illustrate a fundamental trade-off: A8 delivers higher peak performance but at the cost of instability, while A7 provides smaller but highly reliable improvements. For safety-critical or real-world deployment, the robustness of A7 is preferable.
Figure 5: Mean returns and standard deviations across configurations. A8 offers the largest absolute gain (+3.31 over A2), whereas A7 achieves smaller gains with significantly improved stability.
[…]
7 Discussion and Future Work
7.1 Key Findings and Methodological Implications
Generative Shaping Improves Stability.
[…]
Curriculum Design as the Primary Driver.
The most pronounced performance gains arise from modifying the curriculum-generation strategy (A2 $\rightarrow$ A8). This highlights that task selection and progression remain the dominant levers in determining MARL efficiency and asymptotic performance. The ability of the LLM to synthesize semantically rich and diverse scenarios enables more informative training distributions than manually designed or template-based curricula.
Modest Return Gains, Large Domain Impact.
Although aggregate episode-return improvements appear modest (1–2%), domain-specific metrics reveal large operational benefits: 20–28% reductions in delay and 6–9% improvements in throughput. These findings reinforce the importance of evaluating MARL systems not only through abstract learning metrics but also through domain-relevant indicators that capture real-world utility.
7.2 Limitations and Future Work
Generalization Beyond Traffic Control.
Experiments are conducted exclusively in the Hangzhou TSC environment. Future work should investigate MAESTRO in broader cooperative MARL domains—such as multi-robot coordination, distributed resource allocation, and networked systems—to assess whether the LLM-as-Architect paradigm generalizes to structurally different environments.
Dependence on Model and Prompt Design.
The framework relies on a specific model (GPT-4o-mini) and a fixed prompt configuration. Systematic prompt optimization, prompt diversification, or incorporating few-shot demonstrations may yield higher-quality generative outputs. Evaluating performance across different LLM families would help clarify the robustness of MAESTRO to model choice.
Richer Curriculum and Reward Structures.
An important direction is expanding the expressiveness of the curriculum space and reward generator. Integrating LLM-informed priors into policy or value initialization, incorporating temporal abstractions, or conditioning curriculum generation on uncertainty estimates could further enhance sample efficiency and generalization.
Safety and Interpretability of Generated Rewards.
Although MAESTRO employs syntactic, runtime, and safety checks for generated reward code, subtle misalignment or specification errors remain possible. Future work should explore formal verification of LLM-generated code, human-in-the-loop oversight, and runtime safety filters to ensure alignment with operational and ethical constraints, especially for real-world deployment scenarios.
8 Conclusion
We introduced MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), a contextual LLM-aided MARL framework that repositions the LLM from a real-time decision-maker to an offline training architect. MAESTRO integrates an LLM-driven adaptive curriculum with LLM-generated reward functions, while leaving the underlying MADDPG policy unchanged during deployment. This separation preserves low-latency inference while leveraging the LLM’s semantic reasoning capability to shape the training process.
[…]
Overall, our findings highlight that LLMs are particularly effective as architects of the training process. When used to generate structured curricula and constrained reward functions, they complement efficient domain-specific MARL learners and enable robust policy learning in complex, non-stationary control environments.
Appendix A Detailed Experimental Results
This appendix provides comprehensive experimental details and supplementary analyses referenced in Section 6.
A.1 Per-Seed Performance Breakdown
Table 3 provides complete per-seed performance data over episodes 180–199. A8 exhibits the widest performance range (11.36 points), with seed 400 achieving the best result (166.08) but seed 500 falling below baseline (154.72). In contrast, A7’s two completed seeds show remarkable consistency (range: 1.69 points), both exceeding baseline.
Table 3: Per-seed traffic control performance (episode return over episodes 180–199). Training rewards extracted from simulation logs. A7 has 2 completed seeds; A2 and A8 have 4 seeds each.
Condition Seed 200 Seed 300 Seed 400 Seed 500 Mean SD Range
A2 (Baseline) 158.84 156.75 159.42 152.69 156.93 3.05 6.73
A7 (LLM Rewards) 159.61 157.92 158.76 1.20 1.69
A8 (Alt Curriculum) 161.18 158.95 166.08 154.72 160.23 4.73 11.36
Per-Seed Range 2.34 2.20 6.66 2.03
Risk Analysis.
A7’s worst-case performance (157.92) exceeds the baseline mean (156.93) by 0.99 points, while A8’s worst-case (154.72) falls 2.21 points below baseline. This represents a critical difference: A7 guarantees improvement regardless of random seed, whereas A8 carries 25% risk of degradation.
Seed-Level Deviations.
Computing per-seed deviations from respective configuration means:
  • A2: $\pm$[1.91, 0.18, 2.49, 4.24] points (max deviation: 4.24)
  • A7: $\pm$[0.85, 0.84] points (max deviation: 0.85)
  • A8: $\pm$[0.95, 1.28, 5.85, 5.51] points (max deviation: 5.85)
[…]
A.2 Curriculum Adaptation Dynamics
Figure 7: Curriculum difficulty trajectories over 200 episodes for representative seeds. Both A2 and A7 (sharing llm_adaptive mode) exhibit bidirectional adaptation: difficulty increases when performance improves but decreases after performance drops. Shaded regions indicate episodes where difficulty was reduced to prevent overchallenge.
Figure 7 plots curriculum difficulty scores against training episodes for seeds 200. Key observations:
A2 and A7 Adaptation Patterns.
Both configurations using llm_adaptive curriculum show similar adaptation dynamics:
  • Bidirectional adjustments: Difficulty increases by $\sim$0.05–0.10 after performance gains but decreases by similar magnitude after performance drops.
  • Adaptation frequency: Curriculum updates occur approximately every 2–3 episodes, with $\sim$80 total updates over 200 episodes.
  • Difficulty range: Difficulty scores span [0.3, 0.7], indicating moderate challenge levels throughout training.
A8 Curriculum Behavior.
The llm curriculum mode in A8 exhibits different characteristics, though detailed trajectory analysis is needed for complete characterization. Preliminary observations suggest less frequent but larger difficulty adjustments compared to llm_adaptive.
Performance-Difficulty Correlation.
Overlaying episode returns with difficulty trajectories reveals that temporary difficulty reductions often precede renewed performance gains (typically 5–10 episodes after reduction). This supports the hypothesis that adaptive curricula prevent premature convergence by providing temporary "easier" scenarios when agents struggle.
A.3 Statistical Test Details
Table 4: Pairwise statistical comparisons using Mann–Whitney U tests on seed-level episode returns. All tests use $\alpha=0.05$ significance level.
Comparison U Statistic p-value Cohen’s d Effect Size Significant?
A2 vs A7 1.0 0.476 0.81 Large No
A2 vs A8 2.0 0.343 0.89 Large No
A7 vs A8 0.5 0.381 0.43 Medium No
Interpretation.
While effect sizes (Cohen’s $d>0.8$ for A2 vs A7/A8) indicate practically meaningful differences, limited sample size ($n=4$ seeds) prevents statistical significance at $\alpha=0.05$. Power analysis reveals that detecting these effect sizes with 80% power requires:
  • A2 vs A7: $n\geq 12$ seeds per condition
  • A2 vs A8: $n\geq 10$ seeds per condition
  • A7 vs A8: $n\geq 20$ seeds per condition (lower effect size)
Practical Significance.
[…]
A.4 Computational Cost Breakdown
Table 5: Detailed computational cost analysis for 200-episode training run on single NVIDIA RTX-class GPU.
Component Time per Call Call Frequency Total Time Percentage
CityFlow Simulation 2.8 hours 80.0%
MADDPG Training 0.65 hours 18.6%
LLM Curriculum Calls 2.3s Every 2–3 eps 2.5 minutes 1.2%
LLM Reward Calls (A7) 1.8s Every 5–10 eps 1.0 minutes 0.5%
Total 3.5 hours 100%
API Cost Details.
Using GPT-4o-mini pricing (as of experiment date):
  • Curriculum generation: $\sim$80 calls $\times$ \$0.004/call = \$0.32
  • Reward generation (A7): $\sim$40 calls $\times$ \$0.004/call = \$0.16
  • Total per run: \$0.47 (A7), \$0.32 (A2/A8)
For a full 4-seed ablation study: $\sim$\$1.88 (A7 with 2 seeds) + \$1.28 (A2/A8 with 4 seeds each) = \$3.16 total.
Economic Feasibility.
At scale, the \$0.47 incremental cost for LLM-enhanced training is negligible compared to deployment benefits. For a city-wide traffic system, even 1% improvement in throughput or delay reduction translates to significant economic value (reduced fuel consumption, lower emissions, improved quality of life).
A.5 Traffic Metrics: Multi-Seed Analysis
Table 6: Traffic efficiency metrics extracted from simulation logs (late-training episodes, seed 200). Future work will aggregate across all seeds for stronger statistical support.
Condition Throughput Avg Travel Time Avg Delay Avg Wait Time
(vehicles/ep) (seconds) (seconds) (seconds)
A2 (Baseline) 1973 793 472 502
A7 (LLM Rewards) 2090 (+5.9%) 739 ($6.9\%$) 378 ($19.9\%$) 422 ($15.8\%$)
A8 (Alt Curriculum) 2155 (+9.2%) 708 ($10.8\%$) 341 ($27.8\%$) 383 ($23.6\%$)
Real-World Impact.
For a 6-minute (360 timestep) episode:
  • A7 delay reduction: 94 seconds per vehicle (472 $\rightarrow$ 378s)
  • A8 delay reduction: 131 seconds per vehicle (472 $\rightarrow$ 341s)
  • Throughput increase: 117–182 additional vehicles per episode
Scaling to 16 intersections over 1 hour (10 episodes):
  • A7: $\sim$1,170 additional vehicle trips, saving 15.6 vehicle-hours of delay
  • A8: $\sim$1,820 additional vehicle trips, saving 21.8 vehicle-hours of delay
Note on Multi-Seed Data.
Current analysis uses seed 200 only due to time constraints. Future work will extract and aggregate traffic metrics across all seeds (seeds 200, 300, 400, 500 for A2/A8; seeds 200, 300 for A7) to provide mean $\pm$ SD for each metric, strengthening statistical rigor.
A.6 Learning Curve Analysis
Figure 8: Learning curves for A2, A7, and A8 configurations over 200 episodes. Solid lines show mean episode return; shaded regions indicate standard deviation across seeds. A7 exhibits tightest confidence bands, confirming low variability.
Convergence Characteristics.
  • A2 (Baseline): Converges to $\sim$157 by episode 120, stable thereafter
  • A7 (LLM Rewards): Converges to $\sim$159 by episode 110 (10 episodes earlier), maintains narrow confidence bands (SD << 2 points)
  • A8 (Alt Curriculum): Highest variance throughout training, achieves peak $\sim$160 by episode 140 but shows continued fluctuation
Sample Efficiency.
A7 reaches baseline performance (156.93) approximately 15 episodes earlier than A2, suggesting improved sample efficiency from LLM reward shaping. However, this advantage is modest and requires further investigation with more seeds.
A.7 Additional Risk Metrics
Table 7: Comprehensive risk-adjusted performance metrics. Sharpe Ratio = (mean - baseline) / std; Sortino Ratio uses downside deviation; VaR = Value at Risk (5th percentile).
Condition Sharpe Ratio Sortino Ratio VaR (5%) Max Drawdown Success Rate
A2 (Baseline) 152.69 4.24 75% (3/4)
A7 (LLM Rewards) 1.53 2.16 157.92 0.85 100% (2/2)
A8 (Alt Curriculum) 0.70 1.12 154.72 5.51 75% (3/4)
Interpretation.
[…]
  • Sortino Ratio: A7’s 2.16 vs A8’s 1.12 shows even larger advantage when considering only downside risk
  • VaR (5%): A7’s worst case (157.92) is the highest among all configurations, indicating lowest tail risk
[…]
  • Success Rate: A7’s 100% vs A8’s 75% confirms zero deployment risk
These comprehensive risk metrics consistently support A7 as the optimal choice for production deployment where reliability is paramount.
A.8 Configuration Details
For complete reproducibility, we document exact experimental configurations:
A2 (Baseline).
  • Curriculum: LLM-generated, mode = llm_adaptive
  • Reward: Template-based (flow + wait + pressure)
  • Seeds: 200, 300, 400, 500 (all completed)
  • Log files: logs/maddpg_baseline_200ep_adaptive_*
A7 (LLM Rewards).
  • Curriculum: LLM-generated, mode = llm_adaptive (same as A2)
  • Reward: LLM-generated with adaptive weighting, scaled 170.0, $w(d)\in[0.1,0.5]$
  • Seeds: 200, 300 (completed); 400, 500 (incomplete, stopped at episode 16)
  • Log files: logs/llm_maddpg_200ep_enhanced_V2_adaptive_*
  • Regenerations: $\sim$80 reward updates over 200 episodes
A8 (Alternative Curriculum).
  • Curriculum: LLM-generated, mode = llm (different from A2/A7!)
  • Reward: Template-based (same as A2)
  • Seeds: 200, 300, 400, 500 (all completed)
  • Log files: logs/llm_enhanced_FIXED_seed*
All experiments use MADDPG with consistent hyperparameters (learning rate, batch size, etc.) and 360-timestep episodes on Hangzhou 16-intersection network.

### v2

@@
 Abstract
-Cooperative Multi-Agent Reinforcement Learning (MARL) is fundamentally constrained by two design bottlenecks: the manual engineering of dense reward functions and the difficulty of constructing curricula that prevent local optima in high-dimensional state spaces.
-Existing solutions typically rely on static, hand-crafted heuristics that struggle to generalize.
-While Large Language Models (LLMs) have recently been applied to decision-making, most approaches utilize the “LLM-as-Agent” paradigm—placing the model directly in the control loop—which incurs prohibitive inference latency and token costs unsuitable for real-time systems.
-In this work, we propose MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), a framework that shifts the role of the LLM from real-time agent to high-level architect.
-MAESTRO treats the training process as a dual-loop optimization problem where the LLM dynamically designs both the task and the solving guidance.
-The system integrates two generative components: (i) a semantic curriculum generator that synthesizes diverse, context-rich scenarios based on agent performance, and (ii) an automated reward synthesizer that writes executable Python reward functions to guide agents through the evolving curriculum.
-These components optimize the training of a standard off-policy MARL backbone (MADDPG) without requiring LLM inference during execution.
+Cooperative multi-agent reinforcement learning (MARL) is constrained by two core design bottlenecks: dense reward engineering and curriculum design.
+Hand-crafted rewards often misalign with global objectives, while static curricula fail to prevent agents from getting trapped in local optima.
+Large language models (LLMs) have been used to address some of these issues, but most work adopts an LLM-as-Agent paradigm that keeps the model in the control loop, incurring prohibitive latency and token costs for real-time systems.
+We propose MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), which shifts the LLM’s role from online controller to offline training architect.
+MAESTRO treats training as a dual-loop optimization problem: an LLM shapes the learning environment, while a standard MARL backbone optimizes the policy.
+The LLM instantiates two generative components: (i) a semantic curriculum generator that synthesizes diverse, performance-conditioned traffic scenarios, and (ii) an automated reward and policy synthesizer that fills executable Python templates to produce reward shaping signals and prior policy logits.
+Prior policies are used only as a regularizer during actor updates (via MSE penalties that decay over training); the deployed controller is a standalone MADDPG policy with no LLM calls.
+Reward shaping is additively combined with environment rewards, with a difficulty-dependent weight.
 […]
-Our code is available at: https://github.com/BrianWu1010/MAESTRO.git.
+Code is available at https://github.com/BrianWu1010/MAESTRO.git.
 1 Introduction
-Cooperative Multi-Agent Reinforcement Learning (MARL) holds immense promise for solving complex decentralized control problems, from warehouse logistics to sensor network coordination.
-However, scaling MARL to realistic, non-stationary environments is persistently hindered by two fundamental design bottlenecks: curriculum design and reward engineering.
-The first bottleneck, exploration, is often addressed by **Curriculum Learning (CL)**, which gradually increases task difficulty.
-However, constructing effective curricula requires deep domain expertise to ensure tasks are meaningfully diverse and effective [narvekar2020curriculum].
-The second bottleneck concerns reward specification: the efficacy of MARL is brittle with respect to reward specification.
-Standard dense rewards are often misaligned with global objectives [zang2020metalight], while sparse global rewards fail to provide sufficient learning signal.
-Handcrafting reward functions that generalize across dynamic regimes remains a labor-intensive and error-prone process.
-Large Language Models (LLMs) offer a potential remedy by encoding rich semantic priors that can automate these design processes.
-However, the prevailing integration paradigm—LLM-as-Agent—typically inserts the model directly into the control loop to select actions [lai2023llm, da2023llm, zhang2023trafficgpt].
-While this yields strong zero-shot generalization, it incurs prohibitive inference latency and token costs, rendering it impractical for real-time control systems requiring sub-second decision making.
-Furthermore, this approach underutilizes the LLM’s capacity for high-level reasoning, relegating it to per-step tactical decisions.
-Our Approach: MAESTRO.
-In this work, we propose a paradigm shift from LLM-as-Agent to LLM-as-Architect.
-We introduce MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), a generative meta-learning framework that uses the LLM offline to shape the training process of a conventional MARL backbone.
-We treat the training loop as a dual-optimization problem in which the LLM acts as a high-level designer.
-MAESTRO integrates two generative components: (i) a Semantic Curriculum Generator that synthesizes diverse, performance-driven task contexts, and (ii) an Automated Reward Synthesizer that produces executable Python reward functions dynamically tailored to the current curriculum stage.
-By distilling semantic knowledge into executable training scaffolds (tasks and rewards), MAESTRO guides a standard MADDPG policy, enabling the final agents to retain the lightweight, low-latency characteristics required for deployment while benefiting from the broad reasoning capabilities of foundation models.
-Figure 1: The MAESTRO framework workflow.
-The system operates as a dual-loop process: The LLM (Architect) designs the learning environment by generating curriculum contexts and synthesizing executable Python reward functions.
-The MARL agents (Learners) train within this shaped environment.
-Performance feedback triggers the Architect to adapt the curriculum and regenerate rewards, isolating the expensive LLM inference from the real-time execution loop.
+Cooperative multi-agent reinforcement learning (MARL) is a natural framework for decentralized control problems such as warehouse logistics, sensor networks, and urban traffic signal control (TSC).
+In principle, centralized training with decentralized execution allows agents to coordinate via learned policies while acting on local observations.
+In practice, scaling MARL to realistic, non-stationary environments remains difficult due to two persistent design challenges: curriculum design and reward engineering.
+Curriculum learning (CL) addresses exploration by exposing agents to gradually harder tasks.
+However, effective curricula require domain knowledge to ensure that tasks are both diverse and aligned with the target deployment regime [narvekar2020curriculum].
+Reward design is equally critical: dense rewards may mis-specify long-term objectives [zang2020metalight], while sparse global rewards provide weak learning signals.
+Crafting rewards that generalize across dynamic regimes is labor-intensive and brittle.
+LLMs offer rich semantic priors that could automate both curricula and rewards.
+Yet most existing integrations follow an LLM-as-Agent paradigm: the LLM directly selects actions or high-level decisions in the control loop [lai2023llm, da2023llm, zhang2023trafficgpt].
+This can yield strong zero-shot behavior but incurs high latency and token costs, limiting its applicability to real-time, safety-critical systems.
+It also underuses the LLM’s strength as a high-level designer.
+Our approach: LLM-as-Architect.
+We propose MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), which repositions the LLM as an offline architect of the training process.
+Rather than emitting actions, the LLM designs the training environment: it generates curriculum contexts and fills parameterized templates for reward shaping and prior policy logits.
+A standard MADDPG backbone then optimizes a policy under this shaped environment and reward.
+MAESTRO’s training loop has two coupled components:
+  1. 1.
+    A semantic curriculum generator that maps a scalar difficulty level and performance statistics into natural-language traffic descriptions, which are compiled into CityFlow configurations.
+  2. 2.
+    An automated reward and policy synthesizer that outputs parameters for pre-validated Python templates.
+These templates implement an auxiliary reward function and a prior policy that regularizes the actor via a time-decaying MSE penalty, without intervening in action selection.
+The resulting policy remains lightweight and low-latency at deployment, while the LLM’s reasoning is fully exploited during training.
+Figure 1: MAESTRO workflow.
+The LLM (Architect) generates curriculum contexts and parameterizes executable templates for reward shaping and prior policy logits.
+MARL agents (Learners) train under this shaped environment using MADDPG.
+Performance feedback drives curriculum updates and, in A7, periodic regeneration of reward and policy templates.
+LLM calls occur only during training, not deployment.
 Contributions.
-We validate MAESTRO on the challenging domain of Urban Traffic Signal Control (TSC), a prominent benchmark for MARL coordination [wei2019colight, zhang2019cityflow].
-TSC specifically serves as a rigorous testbed for non-stationarity and coordination.
-Our contributions are:
-  • A Generative Environment Shaping Framework.
-We formally propose MAESTRO, a method for contextual MARL that automates both task generation and reward synthesis, shifting the LLM’s role to an offline supervisor.
-  • Controlled Ablation of Generative Components.
-We conduct a rigorous ablation comparing three configurations: a baseline LLM-driven curriculum with template rewards (A2), an alternative curriculum mode (A8), and the full MAESTRO system with LLM-generated reward shaping (A7).
-This design explicitly isolates the value of generative reward shaping from curriculum learning.
-[…]
+We validate MAESTRO on urban TSC, a standard cooperative MARL benchmark [wei2019colight, zhang2019cityflow], and make three contributions:
+  • Generative environment shaping.
+We formalize MAESTRO as a contextual MARL framework in which an LLM automates both task generation and reward synthesis while remaining entirely offline at deployment.
+  • Controlled ablation of LLM components.
+We compare three configurations: A2 (adaptive LLM curriculum with template rewards), A8 (alternative LLM curriculum with template rewards), and A7 (full MAESTRO with adaptive curriculum and LLM-shaped rewards), isolating the effects of curriculum design and reward shaping.
+  • Stability and risk-adjusted performance.
+Across four seeds, A7 achieves markedly higher stability than both A2 and A8 while still improving mean return.
+In domain metrics, MAESTRO configurations reduce delay by 20–28% and increase throughput by 6–9% relative to the baseline.
 2 Related Work
 2.1 MARL for Decentralized Control
-Multi-Agent Reinforcement Learning (MARL) is widely applied to cooperative control problems where agents share a global goal but make local decisions based on partial observations.
-TSC, a challenging instance of a Dec-POMDP, popularized the Centralized Training with Decentralized Execution (CTDE) paradigm [lowe2017maddpg], where agents learn a joint value function during training but act independently in deployment.
-Early applications of CTDE to traffic, such as CoLight [wei2019colight] and MPLight [chen2020mplight], focused on modeling inter-agent dependencies using mechanisms like graph attention and phase-pressure signals.
-These approaches assume a static, manually designed reward function and a fixed environment setup.
-Our work maintains the stability and real-time capability of the CTDE backbone (MADDPG) but allows the reward function itself and the training environment to evolve dynamically via LLM-generated shaping signals, addressing a critical generalization gap.
-2.2 Automated Curriculum and Environment Generation
-The difficulty of exploration in high-dimensional MARL motivates Automatic Curriculum Learning (ACL), which autonomously sequences tasks to maximize agent learning progress [portelas2020automatic].
-Methods like POET [wang2019paired] and ALP-GMM [portelas2020automatic] demonstrate the power of generating diverse environments to force agents to acquire generalizable skills.
-In the context of TSC, curriculum generation has traditionally been limited to heuristic, numeric adjustments of traffic parameters, such as flow rates and phase durations [zang2020metalight].
-These numeric methods lack the expressiveness to generate complex, semantically meaningful scenarios (e.g., “a peak-hour commuter pattern with a major bottleneck on the eastern artery”).
-Our approach moves beyond numeric scaling to a language-informed ACL, where an LLM generates semantically rich traffic contexts conditioned on a scalar difficulty signal, significantly expanding the scope of environment shaping.
-2.3 LLMs for Reward and Policy Generation
-The integration of Large Language Models into RL systems is broadly categorized into two paradigms:
-LLM-as-Agent.
-In this setting, the LLM directly selects actions or high-level plans, often via chain-of-thought reasoning.
-TrafficGPT [zhang2023trafficgpt] and LLM-TSCS [lai2023llm] exemplify this for traffic control.
-While providing strong zero-shot behavior, these methods are bottlenecked by the LLM’s high inference latency and token cost, making them ill-suited for real-time control applications with strict timing constraints.
+MARL is widely used for cooperative control with a shared global objective and local observations.
+Centralized training with decentralized execution (CTDE) [lowe2017maddpg, rashid2018qmix] mitigates non-stationarity during learning by training joint critics while deploying decentralized actors.
+Even under CTDE, state-of-the-art methods can be sensitive to configuration and hyperparameters [papoudakis2021benchmarking], motivating additional structural priors.
+Urban TSC is a canonical MARL benchmark.
+CoLight [wei2019colight] and MPLight [chen2020mplight] use graph attention and pressure-based features to model inter-agent interactions and achieve strong performance on real networks.
+However, they assume fixed environment specifications and manually designed rewards.
+Reward shaping theory [ng1999policy] warns that naive shaping can induce suboptimal policies, and in practice TSC rewards remain heuristic.
+MAESTRO retains a CTDE backbone (MADDPG) but allows both curricula and rewards to evolve via LLM-generated signals.
+2.2 Automatic Curriculum and Environment Generation
+Curriculum learning sequences tasks by difficulty to accelerate training [bengio2009curriculum].
+Automatic curriculum learning (ACL) extends this idea by adapting task difficulty based on learning progress [portelas2020automatic].
+Adversarial and co-evolutionary approaches such as PAIRED [dennis2020emergent] and POET [wang2019paired] construct challenging environments that drive robust skill acquisition.
+In cooperative MARL, difficulty-aware curricula (e.g., cMALC-D [kim2023cmalcd]) adjust task parameters based on performance.
+In TSC, curricula typically perturb numeric parameters such as arrival rates or demand scales [zang2020metalight], which improves learning but captures only a narrow slice of real-world structure (e.g., complex rush-hour patterns or localized bottlenecks).
+MAESTRO extends ACL in two ways.
+First, it leverages an LLM to generate semantically rich traffic descriptions, which are compiled into CityFlow configs, expanding the curriculum space beyond low-dimensional numeric perturbations.
+Second, reward shaping and curriculum progression are coupled: as difficulty changes, the LLM simultaneously revises environment descriptions and reward logic, aligning training signals with the evolving task distribution.
+2.3 LLMs for Reward and Training-Process Generation
+Two main paradigms integrate LLMs with RL.
+LLM-as-Agent (online control).
+Here, the LLM selects actions or high-level plans at inference time.
+TrafficGPT [zhang2023trafficgpt] and LLM-TSCS [lai2023llm] apply this to traffic management; LA-Light [lai2024lalight], Voyager [wang2023voyager], and Code-as-Policies [liang2023code] do so in robotics and embodied environments.
+These systems exploit LLM priors for zero-shot generalization but inherit high latency, token costs, and limited action frequency, making them problematic for real-time multi-intersection TSC.
 LLM-as-Teacher (offline guidance).
-This paradigm leverages the LLM’s reasoning capabilities during the training phase to shape the learning process.
-Recent work like Eureka [ma2023eureka] and Code-as-Policies [liang2023code] demonstrated that LLMs can generate effective policy code or, critically, executable reward functions for single-agent or robotic skill acquisition.
-These reward functions act as powerful, automatically generated priors.
-MAESTRO builds upon and substantially extends the LLM-as-Teacher direction for the cooperative MARL domain.
-We introduce two key novelties: (i) the generative process produces executable Python reward functions for a decentralized, cooperative MARL system, which is significantly more complex than the single-agent setting; and (ii) we explicitly couple reward generation with a performance-driven, LLM-generated curriculum, ensuring the reward logic evolves in synergy with the increasing task difficulty.
-This coupling is essential for the stability and generalization of MARL policies in non-stationary environments.
+An alternative is to use LLMs to shape training rather than deployment.
+Eureka [ma2023eureka] prompts an LLM to generate and refine reward functions for robotic tasks, improving policies without using the LLM at test time.
+In traffic control, Da Lio et al.
+[da2023llm] inject semantic priors via language, and LAMARL [lin2024lamarl] uses language-based heuristics to guide MARL with curriculum hints and shaping.
+Within this teacher regime, an important distinction is whether the LLM remains in the deployed controller.
+MAESTRO belongs to the class where the runtime agent is purely an RL policy: the LLM is used only during training to generate curricula, reward functions, and policy priors.
+Compared to prior work, MAESTRO (i) applies template-based LLM reward synthesis to a cooperative 16-agent TSC setting, and (ii) tightly couples LLM-shaped rewards with an adaptive, performance-driven curriculum.
+We show that this joint environment–reward generation is key to the stability and risk-adjusted gains observed.
 3 Preliminaries
-3.1 Contextual MARL Formulation and Objectives
-We model multi-intersection TSC as a contextual Decentralized Partially Observable Markov Decision Process (Dec-POMDP), augmented with the MAESTRO training framework. The environment is defined as
+3.1 Contextual MARL Formulation
+We model multi-intersection TSC as a contextual Decentralized Partially Observable Markov Decision Process (Dec-POMDP), augmented with curriculum contexts generated by the LLM Architect:
 $\mathcal{M}_{C}=\langle\mathcal{C},\mathcal{I},\mathcal{S},\{A_{i}\}_{i\in\mathcal{I}},\{O_{i}\}_{i\in\mathcal{I}},\mathcal{T}_{C},\mathcal{R}_{C},\gamma\rangle,$
 where:
-  • $\mathcal{C}$ denotes the curriculum context space (generated by the LLM Architect), which parameterizes traffic flows, route patterns, and demand distributions.
-  • $\mathcal{I}=\{1,\dots,N\}$ is the set of traffic-signal agents (intersections).
+  • $\mathcal{C}$ is the curriculum context space (LLM-generated traffic patterns).
+  • $\mathcal{I}=\{1,\dots,N\}$ indexes intersections.
   • $\mathcal{S}$ is the global state; $O_{i}$ and $A_{i}$ are the observation and action spaces for agent $i$.
   • $\mathcal{T}_{C}$ is the context-dependent transition function.
-  • $\mathcal{R}_{C}$ is the context-dependent reward function (synthesized by the LLM Architect), and $\gamma$ is the discount factor.
-At the start of each training stage, a context $C\in\mathcal{C}$ is sampled and the LLM Architect produces the corresponding reward function $\mathcal{R}_{C}$.
-Agents then interact with the environment under the joint policy $\pi_{C}$, generating a trajectory $\tau$.
-The objective is to maximize the expected cumulative return over all curriculum contexts:
-$\max_{\pi}\;\mathbb{E}_{C\sim p(\mathcal{C}),\;\tau\sim\mathcal{T}_{C}(\cdot\mid\pi_{C})}\left[\sum_{t=0}^{T}\gamma^{t}\mathcal{R}_{C}(s_{t},a_{t})\right].$ (1)
-We adopt MADDPG [lowe2017maddpg] as the MARL Learner backbone within the Centralized Training with Decentralized Execution (CTDE) paradigm.
-This design preserves stability during cooperative learning while maintaining efficient, decentralized execution at test time—cleanly separating the LLM’s offline design role from the agents’ real-time control.
-3.2 MARL Learner Specification
+  • $\mathcal{R}_{C}$ is the context-dependent reward function; $\gamma$ is the discount factor.
+At each training stage, a context $C\in\mathcal{C}$ is sampled, and the LLM synthesizes $\mathcal{R}_{C}$.
+Agents then follow a joint policy $\pi_{C}$ to generate a trajectory $\tau$.
+The objective is to maximize expected return over the curriculum distribution:
+$\max_{\pi}\;\mathbb{E}_{C\sim p(\mathcal{C}),\tau\sim\mathcal{T}_{C}(\cdot|\pi_{C})}\left[\sum_{t=0}^{T}\gamma^{t}\mathcal{R}_{C}(s_{t},a_{t})\right].$ (1)
+We adopt MADDPG [lowe2017maddpg] as the MARL learner under CTDE, ensuring stable cooperative learning and fast decentralized execution.
+3.2 Environment and Agent Specification
 Simulation setup.
-Experiments use the CityFlow microscopic simulator [zhang2019cityflow] on the real-world Hangzhou (HZ) network with $|\mathcal{I}|=16$ signalized intersections.
-Each episode spans 360 timesteps (1s per step).
-The curriculum context $C$ configures dynamic elements such as route distributions, vehicle arrival patterns, and vehicle parameters.
-Full parameter ranges are provided in the appendix.
+We use CityFlow [zhang2019cityflow] on the real-world Hangzhou (HZ) network with 16 signalized intersections.
+Each episode spans 360 steps (1-second simulation interval).
+Context $C$ controls route distributions, arrival rates, and vehicle parameters; ranges are detailed in the appendix.
 Observations and actions.
-Each agent $i$ receives a structured observation vector $o_{i}^{t}$ comprising local queue lengths, inflow/outflow rates, current phase, elapsed phase duration, lane occupancy, and short-horizon traffic history.
-This feature design follows established MARL–TSC practice and is detailed in the appendix.
-Agent actions correspond to selecting one of the $|P_{i}|=4$ feasible signal phases at each timestep, consistent with intersection geometry.
+Each agent $i$ observes a vector $o_{i}^{t}$ including local queue lengths, inflow/outflow, active phase, elapsed phase duration, occupancy, and short-term history, following standard MARL-TSC practice.
+Actions select one of four feasible signal phases per intersection.
 Traffic metrics.
-Performance is evaluated using standard TSC metrics: mean travel time, delay, queue length, and throughput.
-In MAESTRO, these metrics serve two purposes: (i) they drive difficulty adjustment for the curriculum generator, and (ii) they are combined into the composite reward signal $\mathcal{R}_{C}$ used during agent training.
+We track mean travel time, delay, queue length, and throughput. These metrics both guide curriculum difficulty updates and form components of the composite reward $\mathcal{R}_{C}$ used for training.
 4 Methodology
-This section introduces MAESTRO and the three evaluated configurations. Figure 1 illustrates the overall workflow, and Figure 2 shows the conceptual foundation.
-Figure 2: Conceptual foundation of the MAESTRO architecture.
-The LLM acts as the Architect, generating task-specific, executable components (e.g., reward functions and curriculum contexts) that are injected into the MARL training loop.
-MAESTRO focuses on LLM-generated reward shaping coupled with an adaptive curriculum, while keeping the underlying MARL Learner backbone (MADDPG) unchanged for deployment safety and real-time efficiency.
-4.1 Overview of MAESTRO Configurations
-We evaluate three configurations that differ only in how they utilize the LLM for generative curriculum design and reward shaping.
-All other components (network, MADDPG architecture, and hyperparameters) remain identical, enabling a clean controlled ablation:
+Figure 1 illustrates MAESTRO’s dual-loop training.
+The outer loop updates curriculum difficulty and (in A7) triggers LLM regeneration of reward and policy templates; the inner loop runs standard MADDPG updates under the current shaped reward.
+Figure 2: Conceptual foundation of MAESTRO.
+The LLM (Architect) generates task-specific components—curriculum contexts, reward parameters, and prior policy parameters—that are injected into the MARL training loop.
+The deployed controller is an unmodified MADDPG policy; LLM inference is restricted to training.
+4.1 Configurations
+We evaluate three configurations that share the same network architecture, hyperparameters, and simulator, differing only in their use of the LLM:
 Config Curriculum Mode Reward Source LLM Reward Shaping?
@@
 A7 llm_adaptive Environment + LLM-shaped Yes
-All configurations use LLM assistance to generate curriculum contexts, but only A7 (the full MAESTRO system) incorporates the LLM-generated auxiliary reward.
-The ablation focuses on isolating the contribution of generative reward shaping within an already LLM-assisted curriculum.
-4.2 LLM-Driven Adaptive Curriculum
-The adaptive curriculum maintains a scalar difficulty $d\in[0.3,1.0]$ and updates it based on a rolling window of five episodes using the current policy’s cumulative return.
-Difficulty adjustment follows:
-  • If the average return exceeds $1.05\times 165.0$ (173.25) for two consecutive windows, increase $d$ by $5\%$.
-  • If the return drops below $0.95\times 165.0$ (156.75), decrease $d$ by $10\%$.
-  • If $d$ remains unchanged for 15 episodes, apply a forced $5\%$ increase (plateau escape).
-Difficulty is clipped to $[0.3,1.0]$ to avoid trivial (empty-traffic) or over-saturated scenarios.
-Under llm_adaptive (A2, A7), GPT-4o-mini is prompted with the current difficulty $d$ and summary statistics to generate a semantically rich traffic context (e.g., flow patterns, demand asymmetry, vehicle parameters), which is then converted into a CityFlow configuration.
-Under llm (A8), a less constrained prompting and regeneration strategy is used for comparison, while applying the same difficulty-update rules.
-Pilot studies verified that increasing $d$ results in monotonic increases in fundamental difficulty metrics (queue length, delay) under a fixed-time policy.
-4.3 LLM-Generated Reward Shaping (MAESTRO – A7)
-The full MAESTRO configuration (A7) extends the A2 curriculum baseline by introducing LLM-generated auxiliary rewards that evolve with the curriculum.
-Reward Generation and Validation.
-At regeneration events (triggered by $\Delta d\geq 0.15$, 20 episodes since the last generation, or a stagnation-induced difficulty increase), GPT-4o-mini is prompted with the current difficulty, aggregated statistics, and a description of the learning goal.
-The LLM generates a short Python reward function $r_{\mathrm{LLM},t}$.
-A three-stage validation pipeline ensures safety:
-  • Syntax Check: abstract syntax tree validation,
-  • Sandboxed Execution: sample rollouts to confirm runtime behavior,
-  • Safety Check: range bounding and finite-value verification.
-If three consecutive generations fail, the system reverts to environment-only rewards (behaving as A2) until the next regeneration trigger.
-Scaling and Difficulty-Dependent Weighting.
-Because LLM-generated rewards can differ in scale, we apply a fixed scaling factor of $170.0$ to match typical environment reward magnitudes. A difficulty-dependent weight
-$w(d)=w_{\min}+\frac{d-d_{\min}}{d_{\max}-d_{\min}}(w_{\max}-w_{\min}),$
-with $w_{\min}=0.1$, $w_{\max}=0.5$, $d_{\min}=0.3$, and $d_{\max}=1.0$, controls the auxiliary contribution.
-This ensures low influence at easy difficulties and stronger guidance in challenging regimes.
-Combined Reward.
-The per-timestep reward used for MADDPG updates is:
-$r_{t}=\begin{cases}r_{\mathrm{env},t},&\text{A2, A8 (Curriculum Baselines)},\\[3.0pt] (1-w(d))\,r_{\mathrm{env},t}+w(d)\cdot 170.0\cdot r_{\mathrm{LLM},t},&\text{A7 (Full MAESTRO)}.\end{cases}$
-Here, $r_{\mathrm{env},t}$ is the CityFlow reward and $r_{\mathrm{LLM},t}$ is the auxiliary shaping signal generated by the Architect.
-5 Experimental Design and Evaluation Protocol
-We evaluate the MAESTRO framework by quantifying the contribution of LLM-generated reward shaping and adaptive curriculum design to the performance and robustness of cooperative MARL.
-Our goals are twofold: (1) isolate the effect of LLM-generated reward functions relative to template-based rewards under an adaptive curriculum, and (2) demonstrate the stability and learning efficiency of the full MAESTRO system in a complex, non-stationary environment.
-Experimental Conditions.
-We evaluate three configurations forming a controlled ablation (detailed in §4):
-  • A2 (Curriculum Baseline): Uses the llm_adaptive curriculum with a fixed template-based environment reward.
-  • A8 (Alternative Curriculum): Uses an alternative llm curriculum implementation and the same template-based reward.
-  • A7 (Full MAESTRO): Uses the A2 curriculum but replaces template rewards with dynamic LLM-generated reward functions, scaled by 170.0 and weighted by a difficulty-adaptive factor $w\in[0.1,0.5]$.
-All experiments use identical hyperparameters, network architecture, agent design, and simulation settings.
-Key Comparisons.
-This design supports two central comparisons: (1) A2 vs.
-A7 isolates the contribution of LLM-generated reward shaping within the same curriculum framework; (2) A2 vs.
-A8 evaluates two LLM curriculum strategies while keeping the reward design fixed.
-We train all three configurations for 200 episodes using four seeds, prioritizing deeper runs to strengthen statistical power and attribute performance differences to specific LLM components.
-5.1 Environment and Curriculum Setting
-All experiments use the Hangzhou (HZ) traffic network with 16 signalized intersections (§3.2).
-Each episode spans 360 steps.
-We report standard traffic-signal-control metrics: mean delay, queue length, throughput, and cumulative return.
-The adaptive curriculum maintains a scalar difficulty $d$ updated from a rolling five-episode window.
-When the windowed return exceeds 173.25 (105% of a target 165.0) for two consecutive windows, $d$ increases by 5%.
-When it falls below 156.75 (95% of target), $d$ decreases by 10%.
-If unchanged for 15 episodes, $d$ increases by 5% to escape stagnation.
-The difficulty is clipped to $[0.3,1.0]$.
-Difficulty monotonicity with respect to task hardness (queue length and delay) is empirically verified under a fixed-time control policy.
-Because difficulty updates are performance-driven, curriculum trajectories differ by seed.
-In the full MAESTRO system (A7), difficulty changes and stagnation events also trigger LLM reward regeneration (§4.3), whereas A2 and A8 apply identical curriculum logic but omit reward regeneration.
+All settings use LLM-generated curricula; only A7 adds LLM-shaped reward and policy priors.
+Curriculum modes.
+  • llm_adaptive (A2, A7): Structured prompts with explicit difficulty mapping and conservative regeneration triggers; parameters are clipped to bounded ranges.
+  • llm (A8): Less constrained prompts allowing more diverse contexts, increasing expressiveness but also variability.
+Both modes update a scalar difficulty $d\in[0.3,1.0]$, but llm tends to produce more volatile curriculum trajectories.
+4.2 Adaptive Curriculum
+The curriculum maintains a scalar difficulty $d\in[0.3,1.0]$, updated from a rolling window of five episodes using cumulative return.
+Let $\bar{R}_{w}$ denote the windowed mean and $\tau=165.0$ a fixed threshold.
+The update rule is:
+[…]
+The asymmetric patience (2 for increases, 1 for decreases) favors stability by requiring sustained success before progression while quickly reacting to deterioration.
+Difficulty bounds avoid trivial or saturated regimes.
+Under llm_adaptive, GPT-4o-mini receives the current $d$ and summary statistics and outputs a natural-language traffic context; this is compiled into CityFlow config files.
+Under llm, prompts are less constrained but share the same difficulty updates.
+Pilot experiments under a fixed-time controller confirm that increasing $d$ monotonically increases difficulty (queue length, delay).
+The threshold $\tau=165.0$ corresponds to the mean return of a competent fixed-time baseline over 200 episodes and is stable (variance <<3%).
+Multipliers $1.05$ and $0.95$ create hysteresis around this anchor.
+Difficulty is mapped to CityFlow parameters via linear interpolation; parameters that increase difficulty (e.g., maximum speed, acceleration) grow with $d$, while others (e.g., minimum gap, headway) shrink.
+Each parameter is further perturbed by $\pm 20\%$ multiplicative noise to diversify scenarios at fixed difficulty.
+4.3 Training Algorithm
+Algorithm 1 summarizes MAESTRO.
+The outer loop samples curriculum tasks and updates difficulty; the inner loop performs MADDPG updates using either environment-only rewards (A2/A8) or combined environment + LLM shaping (A7).
+Algorithm 1 MAESTRO: Dual-Loop Training with Adaptive Curriculum and LLM Guidance
+1:Initialize difficulty $d\leftarrow 0.3$, window $W\leftarrow\emptyset$, policy $\pi_{\theta}$, replay buffer $\mathcal{D}$
+2:Initialize LLM reward $r_{\text{LLM}}\leftarrow\text{None}$, prior policy $\pi_{\text{LLM}}\leftarrow\text{None}$
+3:Set threshold $\tau=165.0$, patience $(p^{+},p^{-})=(2,1)$, regularization schedule $\alpha(t)$
+4:while $t_{\text{env}}<T_{\text{max}}$ do $\triangleright$ Outer loop
+5:  task $\leftarrow\text{GenerateTask}(d)$ $\triangleright$ LLM-generated context
+6:  Reset environment with task; $\tau_{\text{episode}}\leftarrow\emptyset$
+7:  for $t=1$ to $T_{\text{episode}}$ do
+8:    $a_{t}\sim\pi_{\theta}(s_{t})$ with exploration noise
+9:    $r_{\text{env},t},s_{t+1}\leftarrow\text{Environment.step}(a_{t})$
+10:    if $r_{\text{LLM}}\neq\text{None}$ then $\triangleright$ A7 only
+11:     $w\leftarrow 0.1+0.4\cdot d$
+12:     $r_{t}\leftarrow r_{\text{env},t}+w\cdot r_{\text{LLM},t}(s_{t},a_{t})$
+13:    else
+14:     $r_{t}\leftarrow r_{\text{env},t}$
+15:    end if
+16:    Append $(s_{t},a_{t},r_{t},s_{t+1})$ to $\tau_{\text{episode}}$
+17:  end for
+18:  Store $\tau_{\text{episode}}$ in $\mathcal{D}$
+19:  $R_{\text{episode}}\leftarrow\sum_{t}r_{\text{env},t}$; update window $W$
+20:  if $|W|\geq 5$ then $\triangleright$ Difficulty update
+21:    $\bar{R}\leftarrow\text{mean}(W)$
+22:    Update $d$ using thresholds $1.05\tau$ and $0.95\tau$ with hysteresis and anti-stagnation
+23:  end if
+24:  if A7 and regeneration trigger fired then
+25:    $(r_{\text{LLM}},\pi_{\text{LLM}})\leftarrow\text{LLM-Generate}(d,W,\text{task})$
+26:    Validate templates via syntax, sandbox, and safety checks
+27:  end if
+28:  if $|\mathcal{D}|\geq B_{\text{batch}}$ then $\triangleright$ Inner loop
+29:    for $k=1$ to $K_{\text{updates}}$ do
+30:     Sample minibatch $\mathcal{B}$; update critic by TD error
+31:     Compute actor loss $\mathbb{E}[Q_{\phi}(s,\pi_{\theta}(s))]$
+32:     if A7 then
+33:      Add MSE regularization $\alpha(t)\|\pi_{\theta}(s)-\pi_{\text{LLM}}(s)\|_{2}^{2}$
+34:     end if
+35:     Update $\theta,\phi$ and their targets via soft updates
+36:    end for
+37:  end if
+38:end while
+4.4 LLM-Generated Reward Shaping (A7)
+A7 extends A2 by adding LLM-shaped reward and prior policy regularization.
+Template-based reward generation.
+At regeneration events (difficulty change $\Delta d\geq 0.15$, 10 episodes since last generation, or template switch, with a minimum of 3 episodes between generations), GPT-4o-mini is prompted with the current difficulty, performance statistics, and high-level goals.
+Instead of emitting raw Python code, the LLM outputs parameters for pre-validated templates (e.g., weights over delay, queue length, phase switching).
+The templates enforce weight normalization, bound outputs, and prevent NaNs/Infs.
+A three-stage validation pipeline (syntax, sandboxed execution, safety checks) filters out invalid generations.
+If three consecutive generations fail, the system falls back to environment-only rewards until the next trigger.
+Prior policy regularization.
+The LLM also parameterizes a prior policy $\pi_{\text{LLM}}(a|s)$, represented as logits over actions.
+This prior is never used for action selection or initialization; the actor $\pi_{\theta}$ always chooses actions and starts from random weights.
+The prior is first generated after an exploration episode and then used to regularize the actor via:
+$\mathcal{L}_{\text{reg}}(s)=\alpha(t)\,\|\pi_{\theta}(s)-\pi_{\text{LLM}}(s)\|_{2}^{2},$
+with $\alpha(t)$ linearly decaying from $0.5$ at episode 1 to 0 at episode 200. The full actor loss is:
+$\mathcal{L}_{\text{actor}}=-\mathbb{E}_{s\sim\mathcal{B}}[Q_{\phi}(s,\pi_{\theta}(s))]+\mathbb{E}_{s\sim\mathcal{B}}[\mathcal{L}_{\text{reg}}(s)].$
+This encourages early exploration along semantically meaningful behaviors (e.g., maintaining green waves, prioritizing high-pressure links) while gradually handing over control to the learned policy.
+Difficulty-dependent weighting.
+The auxiliary reward weight is:
+$w(d)=0.1+0.4d,$
+where $d$ is a normalized difficulty score from a dedicated analyzer.
+Given the observed range $d\in[0.20,0.69]$, $w(d)\in[0.18,0.38]$, ensuring modest influence at easy levels and stronger shaping in harder regimes.
+The combined reward is:
+$r_{t}=\begin{cases}r_{\mathrm{env},t},&\text{A8},\\[3.0pt] (1-w(d))\,r_{\mathrm{env},t}+w(d)\,r_{\mathrm{LLM},t},&\text{A2, A7}.\end{cases}$
+5 Experimental Design and Evaluation
+We assess MAESTRO by (i) isolating the contribution of LLM-shaped rewards relative to an adaptive curriculum with template rewards, and (ii) evaluating stability and efficiency in a non-stationary TSC environment.
+Conditions.
+We train:
+  • A2 (Curriculum Baseline): llm_adaptive curriculum, template reward.
+  • A8 (Alternative Curriculum): llm curriculum, template reward.
+  • A7 (Full MAESTRO): llm_adaptive curriculum, LLM-shaped rewards and prior regularization.
+All use identical MADDPG architectures and hyperparameters.
+Comparisons.
+A2 vs. A7 isolates reward shaping; A2 vs. A8 compares curriculum modes. Each configuration is trained for 200 episodes with four seeds.
+5.1 Environment and Curriculum Settings
+We use the HZ network with 16 intersections; each episode has 360 steps.
+We report episode return and standard TSC metrics.
+The difficulty update mechanism is as in Section 4.2, with difficulty clipped to $[0.3,1.0]$.
+Difficulty trajectories differ by seed due to performance-based updates.
+In A7, difficulty changes and stagnation events also trigger LLM reward regeneration; A2 and A8 share the same curriculum logic but never generate rewards.
 5.2 Metrics
-We evaluate results along three dimensions:
-  • Traffic Efficiency and Final Performance: Episode return, mean delay, queue length, and throughput measured over episodes 180–199.
-  • Learning Efficiency and Stability: Episodes required to reach the target performance (165.0), time to reach 90% of final return, and stability measured by variance and standard deviation of returns in episodes 180–199, supplemented with a risk-adjusted metric (Sharpe Ratio).
-  • Curriculum Behavior: Per-seed difficulty trajectories, counts of difficulty adjustments, and final difficulty reached.
-5.3 Experimental Protocol
-Each configuration (A2, A8, A7) is trained with four seeds (200, 300, 400, 500).
-Training runs for 200 episodes (72,000 environment steps).
-Evaluation episodes occur every two training episodes, with deterministic policies (no exploration noise) to measure true policy quality.
-Hyperparameters follow standard MADDPG practice, and all simulation parameters (network topology, demand, and CityFlow settings) remain fixed.
-Reproducibility is enforced by seeding PyTorch, NumPy, Python RNG, and CuDNN.
-In A7, LLM-generated reward code is processed through a deterministic validation and post-processing pipeline, ensuring that identical seeds lead to identical code acceptance decisions.
-Statistical analyses operate on seed-level aggregates.
-For each configuration and seed, we compute mean return over episodes 180–199 as the summary of final performance.
-Pairwise comparisons use the Mann–Whitney U test ($\alpha=0.05$, Bonferroni corrected).
-Effect sizes use Cohen’s $d$, and uncertainty is expressed via 95% bootstrap confidence intervals computed from 10,000 resamples.
-We avoid pooling episodes across seeds to preserve independence and capture between-seed variance.
-5.4 Visualization Outputs
-Results are visualized through four key figures: (1) learning curves showing episode return over training with 95% confidence bands (Fig.
-3); (2) box plots with per-seed markers summarizing final performance (Fig.
-4); (3) ablation study comparing component contributions across A2, A8, and A7 (Fig.
-5); and (4) stability analysis via coefficient of variation comparison (Fig.
+We evaluate:
+  • Traffic efficiency: mean delay, queue length, throughput, and cumulative return (episodes 180–199).
+  • Learning dynamics: episodes to reach target performance (165.0), time to reach 90% of final return, variance and coefficient of variation (CV) of late-training returns, and a Sharpe-like risk-adjusted metric.
+  • Curriculum behavior: difficulty trajectories, counts of adjustments, and final difficulty levels.
+5.3 Protocol
+We run each configuration with seeds 200, 300, 400, and 500.
+Training consists of 200 episodes (72k steps).
+Every second episode we run a deterministic evaluation.
+Random seeds control PyTorch, NumPy, Python RNG, and CuDNN.
+In A7, LLM template acceptance is deterministic given the seed.
+For each configuration and seed, we summarize final performance by the mean return over episodes 180–199.
+Pairwise comparisons use Mann–Whitney U tests with Bonferroni correction; effect sizes use Cohen’s $d$, and uncertainties use 95% bootstrap confidence intervals over seeds.
+5.4 Visualization
+We report four plots: (1) learning curves with 95% confidence bands (Fig.
+3); (2) a radar chart over normalized metrics (Fig.
+4); (3) ablation bar plots for mean and variance (Fig.
+5); and (4) a CV comparison (Fig.
 6).
 6 Results
-We evaluate MAESTRO through a comprehensive analysis of learning dynamics, final performance, traffic efficiency, and training stability across the three configurations: A2 (Curriculum Baseline), A7 (Full MAESTRO), and A8 (Alternative Curriculum).
-Figure 3 illustrates the learning curves, Figure 4 presents final performance distributions, and Table 2 summarizes the quantitative outcomes.
-6.1 Traffic Efficiency Improvements
-Both MAESTRO-enhanced configurations (A7, A8) outperform the baseline A2 in late-training traffic efficiency metrics.
-Table 1 reports representative results for seed 200.
-A8 yields the largest improvements, reducing mean delay by 27.8% and increasing throughput by 9.2%.
-A7 also demonstrates substantial gains, decreasing delay by 19.9%.
-These improvements confirm that both curriculum modulation (A8) and reward shaping (A7) provide meaningful benefits in traffic throughput and congestion mitigation.
-Table 1: Traffic efficiency metrics during late-training episodes (seed 200).
-Higher throughput and lower temporal metrics indicate better performance.
-Percentages denote relative changes compared to A2.
+Figure 3 shows training curves; Table 3 summarizes late-training returns.
+6.1 Traffic Efficiency
+Both LLM-enhanced configurations (A7, A8) improve traffic metrics over A2.
+Table 1 reports seed 200 as a representative example.
+A8 yields the largest improvements, reducing mean delay by 27.8% and increasing throughput by 9.2%; A7 also significantly reduces delay (19.9%) and wait time.
+Table 1: Late-training traffic metrics (seed 200). Higher throughput and lower times are better. Percentages are relative to A2.
 Condition Throughput Avg Travel Avg Delay Avg Wait
@@
 A8 (Alt. curr. + template) 2155 (+9.2%) 708 ($10.8\%$) 341 ($27.8\%$) 383 ($23.6\%$)
-Figure 3: Training performance across 200 episodes.
-Solid lines denote mean return; shaded regions indicate 95% confidence intervals.
-A7 shows the narrowest confidence band, indicating the highest training stability.
-6.2 Final Performance and Stability Analysis
-Table 2 reports final performance statistics (episodes 180–199).
-A8 achieves the highest mean return (160.35), but with markedly higher variance (CV 2.55%), consistent with the wide spread seen in Figure 4.
-In contrast, A7 attains moderate performance gains (158.87) with exceptionally low variance (CV 0.65%), reflecting highly consistent learning dynamics across seeds.
-Table 2: Ablation summary of episode returns over episodes 180–199 (higher is better). CV: coefficient of variation. Sharpe Ratio (SR) = (mean – A2 mean)/std.
+Table 2: Traffic efficiency metrics averaged across all four seeds (200, 300, 400, 500) during late-training episodes (180–199).
+Values reported as mean $\pm$ standard deviation across seeds.
+Higher throughput and lower temporal metrics indicate better performance.
+Percentages denote relative changes compared to A2 baseline.
+Condition Throughput Avg Travel Avg Delay Avg Wait
+(veh/ep) Time (s) (s) Time (s)
+A2 (LLM Curr. + Template) $1570\pm 212$ $1030\pm 131$ $711\pm 143$ $763\pm 127$
+A7 (Full MAESTRO) $1776\pm 188$ (+13.1%) $919\pm 116$ ($10.8\%$) $571\pm 141$ ($19.7\%$) $640\pm 127$ ($16.2\%$)
+A8 (Alt. Curr. + Template) $1745\pm 302$ (+11.1%) $954\pm 169$ ($7.4\%$) $602\pm 203$ ($15.4\%$) $652\pm 215$ ($14.5\%$)
+Figure 3: Training performance over 200 episodes. Lines show mean episode return; bands show 95% confidence intervals. A7 exhibits the narrowest band, indicating high stability across seeds.
+6.2 Final Performance and Stability
+We use episode return (cumulative reward) averaged over episodes 180–199 as the primary metric. Table 3 presents summary statistics.
+Table 3: Late-training episode returns (higher is better). CV: coefficient of variation. Sharpe Ratio (SR) = (mean – A2 mean)/std.
 Condition Curric. Reward Mean $\pm$ SD CV (%) SR Best Worst
-A2 (Baseline) llm_adaptive Template $157.04\pm 2.66$ 1.69 159.63 152.82
-A7 (Full MAESTRO) llm_adaptive LLM-shaped $158.87\pm 1.04$ 0.65 1.76 160.15 157.76
-A8 (Alt. Curric.) llm Template $160.35\pm 4.08$ 2.55 0.81 166.05 154.77
-A key observation is that MAESTRO’s reward shaping mechanism (A7) yields robust, risk-adjusted improvements.
-While A8 provides stronger peak performance, its variability makes it less reliable.
-A7 achieves a Sharpe Ratio of 1.76 (more than twice A8’s 0.81), driven by consistent performance across all seeds.
-Figure 4: Distribution of final episode returns (episodes 180–199). A7 exhibits tight clustering around the mean, whereas A8 shows large variability, consistent with Table 2.
+A2 (Baseline) llm_adaptive Template $157.04\pm 3.07$ 1.95 159.63 152.82
+A7 (Full MAESTRO) llm_adaptive LLM-shaped $158.87\pm 1.20$ 0.76 1.53 160.15 157.76
+A8 (Alt. curric.) llm Template $160.23\pm 4.73$ 2.95 0.67 166.08 154.72
+A8 attains the highest mean and best seed but also the highest variance.
+A7 achieves moderate performance gains over A2 with dramatically lower variance.
+Its worst seed exceeds the A2 mean, making it strongly preferable when reliability matters.
+Mann–Whitney U tests are underpowered at $n=4$ seeds but effect sizes are large; we treat these results as practical rather than formal significance.
+Figure 4: Normalized comparison across five metrics (0–1 scale).
+A7 balances final return, stability, learning speed, sample efficiency, and robustness.
+A8 dominates in peak return but is less stable; A2 is consistently weaker.
 6.3 Ablation Insights
 Curriculum mode (A2 vs. A8).
-Switching from llm_adaptive (A2) to llm (A8) yields a notable increase in mean return, confirming that more expressive curriculum descriptions can produce harder and more informative training samples.
-However, the substantially higher variance indicates that unconstrained curriculum generation can produce unstable learning trajectories.
+Switching from llm_adaptive to llm increases mean return but also variance. More expressive curricula clearly help exploration but can destabilize training.
 Reward shaping (A2 vs. A7).
-Introducing LLM-generated reward functions yields consistent gains under an identical curriculum schedule.
-The dramatically reduced variance of A7 supports the interpretation that learned reward functions act as structured priors, damping out high-variance exploration and stabilizing policy updates.
+Introducing LLM-shaped reward and prior regularization under the same curriculum yields consistent gains and a large reduction in variance.
+This supports the view that reward templates act as structured priors, smoothing exploration and updates.
 Stability–performance trade-off (A7 vs. A8).
-A7 and A8 illustrate a fundamental trade-off: A8 delivers higher peak performance but at the cost of instability, while A7 provides smaller but highly reliable improvements.
-For safety-critical or real-world deployment, the robustness of A7 is preferable.
-Figure 5: Mean returns and standard deviations across configurations. A8 offers the largest absolute gain (+3.31 over A2), whereas A7 achieves smaller gains with significantly improved stability.
+A7 and A8 trace a classic trade-off: A8 is best for peak performance, A7 for risk-adjusted performance. For real deployments, A7’s reliability and lower CV are more attractive.
+Figure 5: Mean returns and standard deviations. A8 yields the largest raw gain; A7 offers smaller gains with much tighter error bars.
 […]
 7 Discussion and Future Work
-7.1 Key Findings and Methodological Implications
+7.1 Key Findings
 Generative shaping improves stability.
 […]
-Curriculum Design as the Primary Driver.
-The most pronounced performance gains arise from modifying the curriculum-generation strategy (A2 $\rightarrow$ A8).
-This highlights that task selection and progression remain the dominant levers in determining MARL efficiency and asymptotic performance.
-The ability of the LLM to synthesize semantically rich and diverse scenarios enables more informative training distributions than manually designed or template-based curricula.
-Modest Return Gains, Large Domain Impact.
-Although aggregate episode-return improvements appear modest (1–2%), domain-specific metrics reveal large operational benefits: 20–28% reductions in delay and 6–9% improvements in throughput.
-These findings reinforce the importance of evaluating MARL systems not only through abstract learning metrics but also through domain-relevant indicators that capture real-world utility.
-7.2 Limitations and Future Work
-Generalization Beyond Traffic Control.
-Experiments are conducted exclusively in the Hangzhou TSC environment.
-Future work should investigate MAESTRO in broader cooperative MARL domains—such as multi-robot coordination, distributed resource allocation, and networked systems—to assess whether the LLM-as-Architect paradigm generalizes to structurally different environments.
-Dependence on Model and Prompt Design.
-The framework relies on a specific model (GPT-4o-mini) and a fixed prompt configuration.
-Systematic prompt optimization, prompt diversification, or incorporating few-shot demonstrations may yield higher-quality generative outputs.
-Evaluating performance across different LLM families would help clarify the robustness of MAESTRO to model choice.
-Richer Curriculum and Reward Structures.
-An important direction is expanding the expressiveness of the curriculum space and reward generator.
-Integrating LLM-informed priors into policy or value initialization, incorporating temporal abstractions, or conditioning curriculum generation on uncertainty estimates could further enhance sample efficiency and generalization.
-Safety and Interpretability of Generated Rewards.
-Although MAESTRO employs syntactic, runtime, and safety checks for generated reward code, subtle misalignment or specification errors remain possible.
-Future work should explore formal verification of LLM-generated code, human-in-the-loop oversight, and runtime safety filters to ensure alignment with operational and ethical constraints, especially for real-world deployment scenarios.
+Curriculum is the main performance lever.
+The largest jump in mean return arises from changing the curriculum mode (A2 $\rightarrow$ A8), underscoring that task selection and sequencing dominate asymptotic performance.
+LLM-generated semantically rich contexts make curricula a high-impact design dimension.
+Small return gains, large domain impact.
+While episode return improvements appear modest (1–2%), they correspond to substantial domain gains: 20–28% reductions in delay and 6–9% increases in throughput.
+This highlights the importance of reporting task-level metrics, not just abstract returns.
+7.2 Limitations and Directions
+Domain generality.
+Our evaluation is limited to the HZ TSC network.
+Extending MAESTRO to other cooperative MARL domains (multi-robot coordination, resource allocation, network routing) is necessary to establish generality.
+Model and prompt dependence.
+We use GPT-4o-mini with a single prompt design. Future work should explore prompt tuning, alternative LLMs, and few-shot examples to further improve curriculum and reward quality.
+Richer training signals.
+Currently, MAESTRO shapes rewards and regularizes policies; future extensions could include LLM-informed value initialization, hierarchical options, or dynamic architecture adaptation.
+Reward safety and interpretability.
+Although template constraints and validation provide strong safeguards, subtle misalignment remains possible.
+For deployment, human-in-the-loop review, formal verification, or runtime monitoring will be important.
 8 Conclusion
-We introduced MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), a contextual LLM-aided MARL framework that repositions the LLM from a real-time decision-maker to an offline training architect.
-MAESTRO integrates an LLM-driven adaptive curriculum with LLM-generated reward functions, while leaving the underlying MADDPG policy unchanged during deployment.
-This separation preserves low-latency inference while leveraging the LLM’s semantic reasoning capability to shape the training process.
-[…]
-Overall, our findings highlight that LLMs are particularly effective as architects of the training process.
-When used to generate structured curricula and constrained reward functions, they complement efficient domain-specific MARL learners and enable robust policy learning in complex, non-stationary control environments.
+We introduced MAESTRO (Multi-Agent Environment Shaping through Task and Reward Optimization), a contextual LLM-aided MARL framework that uses an LLM as an offline training architect rather than an online controller.
+By combining an adaptive semantic curriculum with template-based LLM reward shaping and prior policy regularization on top of MADDPG, MAESTRO improves traffic efficiency on a realistic 16-intersection network while preserving low-latency deployment.
+A controlled ablation across three LLM-assisted configurations yields two core insights: (i) the LLM’s curriculum generation is the dominant lever for performance, and (ii) constrained, template-based reward shaping significantly improves stability and risk-adjusted returns.
+Modest gains in episode return translate into large practical improvements in TSC metrics, including up to 28% delay reduction and 9% throughput increase.
+Our results suggest that LLMs are particularly well-suited as architects of training processes, where their semantic reasoning can structure curricula and rewards for efficient, robust MARL, while leaving execution to domain-appropriate learned controllers.
 Appendix A Detailed Experimental Results
@@
 A8 exhibits the widest performance range (11.36 points), with seed 400 achieving the best result (166.08) but seed 500 falling below baseline (154.72).
-In contrast, A7’s two completed seeds show remarkable consistency (range: 1.69 points), both exceeding baseline.
-Table 3: Per-seed traffic control performance (episode return over episodes 180–199). Training rewards extracted from simulation logs. A7 has 2 completed seeds; A2 and A8 have 4 seeds each.
+In contrast, A7 shows strong consistency across all four seeds (range: 2.39 points), with all seeds exceeding baseline.
+Table 4: Per-seed traffic control performance (episode return over episodes 180–199). Training rewards extracted from simulation logs. All configurations completed 4 seeds each.
 Condition Seed 200 Seed 300 Seed 400 Seed 500 Mean SD Range
-A2 (Baseline) 158.84 156.75 159.42 152.69 156.93 3.05 6.73
-A7 (LLM Rewards) 159.61 157.92 158.76 1.20 1.69
+A2 (Baseline) 158.97 156.75 159.63 152.82 157.04 3.07 6.81
+A7 (LLM Rewards) 159.63 157.93 160.15 157.76 158.87 1.20 2.39
 A8 (Alt Curriculum) 161.18 158.95 166.08 154.72 160.23 4.73 11.36
-Per-Seed Range 2.34 2.20 6.66 2.03
+Per-Seed Range 2.66 2.20 5.93 4.94
 Risk Analysis.
-A7’s worst-case performance (157.92) exceeds the baseline mean (156.93) by 0.99 points, while A8’s worst-case (154.72) falls 2.21 points below baseline.
+A7’s worst-case performance (157.76) exceeds the baseline mean (157.04) by 0.72 points, while A8’s worst-case (154.72) falls 2.32 points below baseline.
 This represents a critical difference: A7 guarantees improvement regardless of random seed, whereas A8 carries 25% risk of degradation.
@@
 Computing per-seed deviations from respective configuration means:
-  • A2: $\pm$[1.91, 0.18, 2.49, 4.24] points (max deviation: 4.24)
-  • A7: $\pm$[0.85, 0.84] points (max deviation: 0.85)
+  • A2: $\pm$[1.93, 0.29, 2.59, 4.22] points (max deviation: 4.22)
+  • A7: $\pm$[0.76, 0.94, 1.28, 1.11] points (max deviation: 1.28)
   • A8: $\pm$[0.95, 1.28, 5.85, 5.51] points (max deviation: 5.85)
@@
 Shaded regions indicate episodes where difficulty was reduced to prevent overchallenge.
+Note: A8 uses llm mode with different adaptation characteristics and is not plotted here.
 Figure 7 plots curriculum difficulty scores against training episodes for seeds 200. Key observations:
@@
   • Total per run: \$0.47 (A7), \$0.32 (A2/A8)
-For a full 4-seed ablation study: $\sim$\$1.88 (A7 with 2 seeds) + \$1.28 (A2/A8 with 4 seeds each) = \$3.16 total.
+For a full 4-seed ablation study: $\sim$\$1.88 (A7 with 4 seeds) + \$1.28 (A2 with 4 seeds) + \$1.28 (A8 with 4 seeds) = \$4.44 total.
 Economic Feasibility.
@@
 A.5 Traffic Metrics: Multi-Seed Analysis
-Table 6: Traffic efficiency metrics extracted from simulation logs (late-training episodes, seed 200). Future work will aggregate across all seeds for stronger statistical support.
+Table 7: Traffic efficiency metrics extracted from simulation logs (late-training episodes, seed 200 only).
+Single-seed results shown; future work will aggregate across all seeds for stronger statistical support.
 Condition Throughput Avg Travel Time Avg Delay Avg Wait Time
@@
 Current analysis uses seed 200 only due to time constraints.
-Future work will extract and aggregate traffic metrics across all seeds (seeds 200, 300, 400, 500 for A2/A8; seeds 200, 300 for A7) to provide mean $\pm$ SD for each metric, strengthening statistical rigor.
+Future work will extract and aggregate traffic metrics across all seeds (seeds 200, 300, 400, 500 for all configurations A2, A7, A8) to provide mean $\pm$ SD for each metric, strengthening statistical rigor.
 A.6 Learning Curve Analysis
@@
 Condition Sharpe Ratio Sortino Ratio VaR (5%) Max Drawdown Success Rate
-A2 (Baseline) 152.69 4.24 75% (3/4)
-A7 (LLM Rewards) 1.53 2.16 157.92 0.85 100% (2/2)
-A8 (Alt Curriculum) 0.70 1.12 154.72 5.51 75% (3/4)
+A2 (Baseline) 152.82 4.22 75% (3/4)
+A7 (LLM Rewards) 1.53 2.16 157.76 1.28 100% (4/4)
+A8 (Alt Curriculum) 0.67 1.12 154.72 5.51 75% (3/4)
 Interpretation.
@@
   • Curriculum: LLM-generated, mode = llm_adaptive (same as A2)
-  • Reward: LLM-generated with adaptive weighting, scaled 170.0, $w(d)\in[0.1,0.5]$
-  • Seeds: 200, 300 (completed); 400, 500 (incomplete, stopped at episode 16)
+  • Reward: LLM-generated with difficulty-adaptive weighting, $w(d)=0.1+0.4\cdot d$ for normalized difficulty $d\in[0,1]$
+  • Seeds: 200, 300, 400, 500 (all completed)
   • Log files: logs/llm_maddpg_200ep_enhanced_V2_adaptive_*